Find the volume of solid of revolution generated when the area under a curve y*2=x+2 is rotated about the x-axis , between x=0 and x=4.
1
Expert's answer
2019-11-06T01:24:10-0500

Solution To find the volume, we use the formula


"V=\pi \int_{a}^b y^2dx"

where y(x) is function is rotated about the x-axis , between x=a and x=b.

For our case we get y=(x+2)/2; a=0 and b=4. Therefore


"V=\pi \int_{0}^4 (\frac {x+2} {2})^2dx=\pi \frac {(x+2)^3}{12} |_0^4=\pi (18-\frac{2}{3})=\frac {52} {3}\pi"



Answer

"V=\frac {52} {3}\pi"